# TextOCVP: Object-Centric Video Prediction with Language Guidance

### v2

@@ 3.4 Training and Inference
Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
Our predictor is trained by minimizing the following loss:
@@ 4.1 Experimental Setup
CATER (Girdhar and Ramanan 2020) is a dataset of long video sequences showing 3D objects in motion, each accompanied by a descriptive caption. We used the CATER-hard variant (Hu, Luo, and Chen 2022), containing 30,000 sequences with $64\times 64$ frames featuring between two and eight objects, two of which follow the scripted motion.
CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset featuring video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where the caption describes placing a block into a bowl.
Baselines
@@
Refer to captiont=1t=122101020203030MAGE TextOCVP
‘the medium green metal sphere is sliding
to (2, 1). the small brown metal cube
is picked up and placed to (-3, 1)’
Figure 4: Text-guided video prediction result on CATER. Top row shows ground truth. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
Implementation Details
All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for scene parsing and video rendering. $\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al. 2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al. 2019) to decode the slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images. On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
Further experiment details are provided in Appendix B-E.
@@ 4.2 Results
CATER Results
Fig. 4 qualitatively compares TextOCVP and MAGE. Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics.
[…]
(a) CATER Evaluation
@@
OCVP 0.874 0.078 4.16 0.854 0.101 8.08
Non-OC 0.874 0.092 3.04 0.849 0.112 8.62
SEER 0.723 0.245 11.23 0.535 0.299 17.29
MAGE 0.877 0.108 3.46 0.871 0.111 5.88
TextOCVP 0.922 0.036 2.16 0.902 0.044 5.09
(b) CLIPort Evaluation
@@
Non-OC 0.901 0.184 8.13 0.872 0.210 13.23
SEER 0.887 0.141 6.80 0.622 0.331 8.29
$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 2.11 0.931 0.075 2.59
TextOCVP 0.950 0.062 1.36 0.931 0.078 2.23
Table 2: Ablation studies. We ablate different architectural design choices in our proposed TextOCVP model.
(a) Effect of number of layers in predictor.
@@
(b) Evaluation of different text encoders on CATER and CLIPort.
@@
No 0.946 0.066 1.58
Yes 0.950 0.062 1.36
(d) Effect of the number of slots $N_{\mathbf{S}}$.
@@
CLIPort Results
‘put the blue block in the gray bowl
[…]
TextOCVPt=1t=12020303040405050Missing
[…]
Nevertheless, we identify that TextOCVP often generates visual artifacts and lacks fine-grained textures, which can affect its quantitative scores despite accurate predictions.
Further figures and results are provided in Appendix F.
4.3 Model Analysis
Ablation Studies
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance. The results are presented in Table 2(d).
Number of Layers (Table 2(a))   TextOCVP performs best with $N_{\text{P}}=8$ predictor layers, outperforming shallower variants. We did not explore scaling beyond eight layers due to the significant increase in the number of parameters.
Residual Connection (Table 2(c))   Incorporating a residual connection in the predictor, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
Text Encoder (Table 2(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5. A frozen T5 module led to the best TextOCVP performance on the CLIPort dataset. Although all encoders led to similar performance on the CATER dataset, we opted for the frozen T5 to avoid unnecessarily increasing the model’s parameter count.
@@ 4.3 Model Analysis
Number of Slots (Table 2(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background). This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al. 2024).
@@
Table 3: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video prediction performance and its drop relative to the known-color dataset.
@@
$\text{MAGE}_{\text{DINO}}$ 0.935 -0.5% 0.076 -19% 0.924 -0.7% 0.087 -16%
TextOCVP 0.946-0.4% 0.066-6.4% 0.927-0.4% 0.083-6.4%
Model Robustness
‘the large purple rubber cone is picked up
and placed to (2, 3). the small gold metal
snitch is picked up and placed to (-1, 1).’
Original
Caption
a)
‘the medium cyan rubber sphere is picked up
and placed to (-2, 2). the medium purple
metal cone is sliding to (2, 3).’
Changed
Objects
& Actions
b)
‘the large purple rubber cone is sliding to
(-1, 1). the small gold metal snitch is
rotating. the medium purple metal cone
is picked up and placed to (-1, -3).’
More
Moving
Objects
c)
t=1t=12215153030
Figure 6: Evaluation of TextOCVP’s controllability on CATER. Top row shows the ground truth sequence. TextOCVP enables predicting different sequence continuations conditioned on distinct text instructions.
We evaluate the generalization ability of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training. This setup assesses the models’ ability to generalize to novel scene configurations and adapt to unseen text-visual combinations.
Table 3 presents the results on this evaluation set with unseen colors. We report both the video prediction performance as well as the performance drop relative to the evaluation set with known colors. Our model consistently outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons. Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric. These results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction—particularly in contrast to holistic scene representations that struggle with novel scene compositions.
Controllability
A key objective of text-driven video prediction is to provide fine-grained control over the prediction process, which is enabled through language instructions describing the objects in the scene and their expected motion.
In Fig. 6, we qualitatively showcase the level of control that TextOCVP provides over its generations on the CATER dataset. We qualitatively evaluate this ability by generating multiple video continuations from the same reference frame while varying the language instruction. These variations include changes to both the target objects and their actions, as well as providing instructions that include a number of actions beyond what the model was trained on. As shown in Fig. 6, TextOCVP successfully identifies the relevant objects and executes the actions described in the text instructions. Notably, our model distinguishes between two nearly identical purple cones, despite the identical shapes and color, and correctly generates a sequence with the specified motion.
These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism, enabling the predictor to accurately forecast the motion of each individual object. This highlights the benefits of combining object-centric representations with text guidance for fine-grained, controllable video prediction.
Interpretability
Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption2020Refer to caption3030
‘the cone is sliding to (-3, -3).
the snitch is picked up and placed to (-1, 2 ).’
Refer to captionRefer to caption1.00.50.0Refer to caption

### v3

@@ 3.4 Training and Inference
Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, the predictor forecasts subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time. TextOCVP predictor is trained by minimizing the following loss:
@@ 4.1.1 Datasets
CATER (Girdhar and Ramanan, 2020) is a dataset of long video sequences showing 3D objects in motion, each paired with a descriptive caption. We used the CATER-hard variant (Hu et al., 2022), containing 30,000 sequences with $64\times 64$ frames featuring two to eight objects, two of which follow the scripted motion.
CLIPort (Shridhar et al., 2022) is a robot manipulation dataset featuring video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where each scene contains six objects, either a block or a bowl, on a 2D table plane, and a robot arm. The caption describes placing a block into a bowl.
4.1.2 Baselines
@@
4.1.3 Implementation Details
All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for scene parsing and video rendering. $\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al., 2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al., 2019) to decode slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images. On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024. Further experimental details—including datasets, baselines, evaluation metrics, and implementation details—are provided in Appendices B-E.
@@ 4.2 Results
4.2.1 CATER Results
Table 1: Quantitative evaluation on CATER and CLIPort datasets for prediction horizons of $T=9$ and $T=19$. TextOCVP outperforms the baselines. Best two results are shown in bold and underlined, respectively.
(a) Quantitative evaluation on CATER.
@@ 4.2.1 CATER Results
OCVP (Villar-Corrales et al., 2023) 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62
SEER (Gu et al., 2024) 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29
MAGE (Hu et al., 2022) 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88
TextOCVP (Ours) 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09
(b) Quantitative evaluation on CLIPort.
@@
Non-OC 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
SEER (Gu et al., 2024) 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
TextOCVP (Ours) 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
Fig. 4(a) qualitatively compares TextOCVP and MAGE on CATER. TextOCVP generates a sequence that exhibits coherent object trajectories and accurate motion grounded in the input textual instruction. This alignment is made possible by the explicit object-centric design, which disentangles the scene into structured object representations and allows the model to reason about each object’s dynamics independently. In contrast, MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. These issues arise from the difficulty of modeling object-specific dynamics in pixel or token space without explicit object-level abstraction.
These results demonstrate the importance of structured, object-centric representations in capturing the compositional and controllable nature of physical scenes, which is especially critical in tasks involving fine-grained object manipulations.
Refer to captiont=1t=122101020203030MAGETextOCVP‘the medium green metal sphere is sliding to (2, 1).the small brown metal cube is picked up and placed to (-3, 1)BlurredChanged
(a) Text-guided video prediction result on CATER. Top row depicts the ground truth frames. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
[…]
(b) Qualitative evaluation on CLIPort for text-driven video prediction. Top row depicts the ground truth frames. TextOCVP successfully completes the pick-and-place task, whereas in $\text{MAGE}_{\text{DINO}}$ the moved block disappears.
Figure 4: Qualitative comparison of TextOCVP and baseline methods on text-guided video prediction.
4.2.2 CLIPort Results
[…]
Nevertheless, we observe that TextOCVP often generates visual artifacts and lacks fine-grained textures. While these imperfections can affect its quantitative scores, TextOCVP’s object-centric structure enables robust long-term prediction and precise instruction-following behavior.
4.3 Model Analysis
4.3.1 Ablation Studies
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance. The results are presented in Table 2, and analysed below across four main design axes. Further results and analysis are provided in Appendix F.
Table 2: Ablation studies on key design choices in TextOCVP, including predictor depth (2(a)), text encoder choice (2(b)), residual connections (2(c)), and slot count (2(d)).
(a) Effect of number of layers ($N_{\text{P}}$) in predictor module.
@@ 4.3.1 Ablation Studies
(b) Impact of different text encoders, including a transformer encoder (TF) and two T5 variants, on CATER and CLIPort.
@@
0.946 0.066 1.58
0.950 0.062 1.36
(d) Effect of the number of object slots ($N_{\mathbf{S}}$).
@@
Number of Layers (Table 2(a))
We study the effect of increasing the number of transformer layers ($N_{\text{P}}$) in the predictor module. We observe that prediction quality improves with model depth, with $N_{\text{P}}=8$ yielding the best performance across all metrics on the CATER dataset. This suggests that deeper predictor models enable more accurate modeling of object dynamics. However, we did not explore beyond eight layers due to the substantial increase in computational cost, training time and parameter count.
Residual Connection (Table 2(c))
We evaluate the effect of introducing a residual connection in the predictor module, where the updated slot representation is defined as $\hat{\textbf{S}}_{t+1}=\textbf{S}_{t}+f_{\text{pred}}(\textbf{S}_{t})$. Adding this residual path improves performance across all metrics on the CLIPort dataset, particularly in improving the temporal consistency of predicted slots. This result is consistent with prior findings (Villar-Corrales et al., 2023), which show that residual updates are beneficial for iterative refinement in structured latent spaces.
Text Encoder (Table 2(b))
We evaluate the performance of three different text encoders, including a lightweight transformer trained from scratch (TF), a frozen T5 encoder, and a fine-tuned (FT) T5. While all three perform similarly on CATER, the frozen T5 text encoder achieves the best results on CLIPort, especially on perceptual metrics such as JEDi and LPIPS. Interestingly, fine-tuning the T5 encoder led to slightly degraded performance, likely due to overfitting or interference with the pretrained representations. We therefore use the frozen T5 encoder in all main experiments to balance performance and model efficiency.
Number of Slots (Table 2(d))
@@
Finally, we explore the impact of varying the number of slots ($N_{\mathbf{S}}$) used to represent the scene. Although each CLIPort scene can be described with eight slots—representing six objects, the robot arm, and the background—we find that using ten slots results in notably better performance. This observation suggests that the additional slots can function as internal registers, supporting attention routing or acting as a form of cache that aids with internal model computations (Darcet et al., 2024).
4.3.2 Model Robustness and Generalization
We evaluate the robustness and generalization of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on two CLIPort evaluation sets: one involving color variations in the text instructions that were not encountered during training (unseen-color), and another featuring scenes with more objects than observed during training, eight instead of six (more-objects). These experiments assess model robustness under distribution shifts and test the ability to generalize to unseen visual-linguistic combinations and novel scene configurations. Results are summarized in Table 3, reporting both the absolute performance and the relative drop compared to the standard (seen-color or six-object) settings.
On the unseen-color benchmark (Table 3(a)), TextOCVP consistently outperforms $\text{MAGE}_{\text{DINO}}$ across all metrics and prediction horizons, showing significantly smaller degradation in perceptual similarity (LPIPS). This indicates stronger robustness to novel color-object combinations. In the more-objects setting (Table 3(b)), TextOCVP again maintains superior performance with a notably smaller decline in LPIPS, demonstrating resilience to higher scene complexity and unseen object counts. Qualitative examples in Appendix F.6 further support these findings.
Overall, these results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction tasks, particularly in contrast to holistic scene representations that struggle with novel scene compositions.
@@ 4.3.2 Model Robustness and Generalization
Table 3: Quantitative evaluation on two CLIPort test sets with unseen colors and larger number of objects. We report the absolute video prediction performance and the relative drop (in parentheses) compared to the original evaluation setting. Best result is marked in bold.
(a) Evaluation on CLIPort test set with objects of colors unseen during training.
@@
$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 0.935 (-0.5%) 0.076 (-19%) 0.924 (-0.7%) 0.087 (-16%)
TextOCVP (Ours) 0.946(-0.4%) 0.066(-6.4%) 0.927(-0.4%) 0.083(-6.4%)
(b) Evaluation on CLIPort test set with a larger object count than training scenes.
CLIPort1→9 CLIPort1→19
[…]
$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 0.929 (-1.2%) 0.088 (-37.5%) 0.920 (-1.2%) 0.094 (-25.3%)
TextOCVP (Ours) 0.936 (-1.5%) 0.076(-22.6%) 0.921(-1.1%) 0.090(-15.4%)
4.3.3 Interpretability
Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption1515Refer to caption2020Refer to caption2525Refer to caption3030‘the cone is sliding to (-3, -3). the snitch is picked up and placed to (-1, 2).’ Refer to captionRefer to caption1.00.50.0Refer to caption